# Top row: MNIST digit 5 subject to 10 increasing levels of pixelation Bottom row: Average softmax for all MNIST digit 5 classes across 12 perturbation types, subject to 10 increasing perturbation levels

### v2

@@ 1 Introduction
Distribution shift occurs when the training data distribution differs from the deployment data distribution, leading to a decline in model performance [quinonero2009dataset, hendrycks2019benchmarking]. It can manifest as covariate shift, concept drift, or domain shift [moreno2012unifying, shimodaira2000improving, gama2014survey, patel2015visual]. To quantify and mitigate distribution shift, researchers employ statistical divergence measures [kullback1951information, gretton2012kernel], domain adaptation techniques [wang2018deep, sugiyama2007covariate, pan2009survey, ganin2016domain], and adaptive algorithms [lu2018learning, baena2006early].
@@ 2.1 Softmax Output And Environmental Aspects
Furthermore, there is a growing awareness of the need for sustainable AI practices and the development of green AI frameworks [schwartz2020green]. These initiatives focus on designing and implementing machine learning systems that prioritize energy efficiency, resource conservation, and environmental sustainability. By adopting sustainable AI practices, we can mitigate the environmental impact of discarded softmax outputs and other inefficiencies in machine learning pipelines.
@@ 2.2 Using All Softmax Output Predictions
Utilizing the entire set of softmax prediction probabilities, instead of solely depending on the maximum output, has been widely investigated to improve the safety, robustness, and trustworthiness of machine learning models. By considering the complete distribution of class predictions, more reliable and informative prediction pipelines can be developed, enabling techniques such as uncertainty quantification, anomaly detection, and robustness against adversarial attacks [kendall2017uncertainties, lakshminarayanan2017simple, hendrycks17baseline, goodfellow2014explaining, szegedy2013intriguing, madry2017towards, guo2018countering].
@@
The Misclassification Likelihood Matrix (MLMs) proposed in this study we believe has potential to further enhance the interpretability and risk mitigation capabilities of softmax probabilities. MLMs provide a comprehensive view of the model’s misclassification tendencies by capturing the likelihood of each class being misclassified as another class. By analyzing the patterns and magnitudes of these misclassification likelihoods, decision-makers can identify the most common and critical sources of errors. This information can be used to prioritize model improvements, guide data collection efforts, and establish decision thresholds based on the acceptable level of risk. Moreover, MLMs can be employed to generate explanations for the model’s predictions, highlighting the classes that are most likely to be confused and the factors contributing to the misclassifications. This transparency enables users to better understand the limitations and potential failure modes of the model, promoting informed decision-making and fostering trust in the system.
@@ 3.1 Neural Networks
Figure 1: MNIST testing dataset confusion matrix where then least misclassified digit is 0 (top row) and the most misclassified digit is 8 (second row from bottom up).
@@
We study not only the highest probability score (prediction), but the entire neural network output vector $\mathbf{p}=(p_{1},p_{2},\dots,p_{K})$ where $\sum p_{i}=1$, representing a probability distribution obtained by normalizing the logit vector $\mathbf{z}=(z_{1},z_{2},\dots,z_{K})$ through the softmax function, $p_{i}=\text{softmax}(z_{i})=e^{z_{i}}/\sum_{j=1}^{K}e^{z_{j}}$. For example, given $\mathbf{p}=[0.01,0.01,0.01,0.01,0.9,0.01,0.01,0.01,0.01,0.01]$, the predicted class is ’4’, corresponding to the highest value at index five, reflecting the confidence of the prediction for each class from ’0’ to ’9’. The logits, representing log-likelihoods of class memberships, are related to probabilities by $z_{i}=\log(p_{i}/(1-p_{i}))$ where $z_{i}$ is the logit for class $i$, and $p_{i}$ is the probability of the input belonging to class $i$[goodfellow2016deep, bishop2006pattern].
@@
The model was trained using the Stochastic Gradient Descent (SGD) optimizer with a learning rate of 0.01 and a batch size of 64. The learning rate was determined using a custom learning rate function that decreases the learning rate over time. The Cross-Entropy Loss function was used as the criterion for optimization.
@@ 3.2 Perturbed MNIST Dataset
To test our trained model, we augment the MNIST testing dataset, consisting of 10,000 images with 10 classes of approximately 1,000 examples each. We create 12 perturbations - Brightness, Contrast, Defocus Blur, Fog, Frost, Gaussian Noise, Impulse Noise, Motion Blur, Pixelation, Shot Noise, Snow and Zoom Blur - following prior work by [hendrycks2019benchmarking], at 10 intensity levels, generating a dataset of 1,2010,000 images. We modify perturbation function parameters by trial and error such that increased perturbation levels on the testing dataset cause network predictive accuracies to decrease linearly.
@@ 3.3 Misclassification Likelihood Matrix
Figure 2: Perturbed MNIST testing dataset MLM, where dark blue indicates greater, and lighter yellow lower likelihood of misclassification.
@@
After obtaining the initial centroids, we use them to initialize the KMeans algorithm. The KMeans algorithm iteratively assigns each softmax output to the nearest centroid and updates the centroids based on the assigned outputs until convergence. This process refines the centroids and yields the final adjusted centroids $\hat{\mu}_{0},\hat{\mu}_{1},...,\hat{\mu}_{9}$.
b. Obtaining the Distance to a Neighboring Centroid: Given a test example $x$ with true class label $y$, we obtain the softmax output $s=f_{\theta}(x)$. To calculate the distance between $s$ and a neighboring class centroid $\hat{\mu}c$, where $c\neq y$, we use the Euclidean distance:
@@ 3.4 Testing the hypothesis
Let $L^{(p)}\in\mathbb{R}^{10\times 10}$ be the misclassification likelihood matrix at perturbation level $p$, where $p\in\{0,1,...,9\}$. Each element $L^{(p)}_{y,c}$ represents the likelihood of an example from class $y$ being misclassified as class $c$ at perturbation level $p$.
@@
$\sigma_{y,c}=\sqrt{\frac{1}{10}\sum_{p=0}^{9}(L^{(p)}_{y,c}-\bar{L}_{y,c})^{2}}$ (8)
@@ 4.2 Clustering results
Table 1 presents the Euclidean distances between the average softmax outputs and the K-Means centroids for each digit class (0-9). For most digit classes (0, 1, 2, 3, 4, 7, 9), the Euclidean distances are extremely small, on the order of 1e-14 to 1e-15, indicating that the average softmax outputs are very close to the centroids obtained from the K-Means clustering algorithm. However, for digit classes 5, 6, and 8, the Euclidean distances are slightly larger, on the order of 1e-4, suggesting a minor difference between the average softmax outputs and the K-Means centroids for these specific classes, which could be attributed to the presence of some misclassified or ambiguous examples within these classes. Despite the slight deviations observed for classes 5, 6, and 8, the overall distances between the average softmax outputs and the K-Means centroids are relatively small across all digit classes, indicating a strong similarity between the representations learned by the neural network and the centroids obtained from the K-Means clustering algorithm. The small distances suggest that using the average softmax outputs as initial centroids for the K-Means clustering algorithm is an effective approach, as they provide a good starting point for the clustering process and can potentially lead to faster convergence.
@@ 4.3 Misclustered Images
Looking at the pattern of the softmax output (network prediction) gives a good indication of how confident a model is of the class prediction. Figure 3 contains two bar charts that show the MNIST-trained CNN softmax output for MNIST dataset digit ids 8688 and 22561 (starting from index 0), where both are labelled as class 6 and correctly classified by the CNN as such. The bar chart on the left shows the softmax output assigns a very high probability to both classes 5 and 6 (where 6 is the highest). The bar chart on the right shows that the CNN assigns a high probability to classes 6 and 8 (where 6 is again the highest). The cluster labels however do not correspond as digit id 8688 is nearer to the cluster 5 centroid, and digit id 22561 is nearer to cluster 8 centroid. This volatility may be explained by the relatively high deltas shown in Table 1, where the K-Means centroids moved a relatively large distance with respect to the initial values (given by average softmax outputs for correct digit classifications in the training dataset), for digit classes 5, 6 and 8.
Conversely, the pattern of the distance to class centroids for a given softmax output is a good indication of how close the softmax output is to a class centroid with respect to other class centroids. Figure 4 shows on the left the distance to all class centroids of digit id 8688 network softmax output, placing it closing to centroids 5 and 6, while the image on the right (digit id 22561) places the softmax output closer to centroids 6 and 9. We notice the inverse relation between both sets of figures, where the highest bars (network confidence) in the first set correspond to the lowest bars (distance to class centroid) set in the second set.
@@ 4.4 Applying Perturbations
Through trial and error, we find sets of parameters values for each perturbation function, such that as the level increases from 1 to 10, the network predictive accuracy decreases from 98.52% maximum to 36.32% minimum and from average 97.61% (level 1) to 48.83% (level 10). The accuracy degradation can be observed both Figures 5 and 6.
@@
Figure 7 shows a heatmap that visualizes the accuracy trained on the original MNIST training dataset and tested with our perturbed dataset described in section 3.2.
@@ 4.5 Misclassification Likelihood Matrix
Figure 2 shows the Misclassification Likelihood Matrix obtained by the method described in section 3.3. Comparing the figure with the MNIST testing (non-perturbed) dataset shown in Figure 1 we can notice some information added by the MLM that may help inform risk assessment. On the top right of the confusion matrix we seea cluster of zeros, representing digits 0, 1 and 2 misclassifications of 8 and 9. The MLM is able to provide fine-grained information about the misclassification likelihood, assigning relatively low albeit non-zero likelihood misclassification values. As the MLM represents a much larger dataset, there are some noticeable discrepancies. The MLM does not assign a high likelihood for digit 9 misclassified as 7, which is the highest in the testing dataset. On the other hand, row 7 is a close match. with 7 being misclassified as 2 holds in the perturbed MNIST dataset.
@@ 5 Conclusions and Future Work
In this study, we proposed a novel approach for quantifying the reliability of neural network predictions under distribution shifts by leveraging clustering techniques and analyzing the distances between softmax outputs and class centroids. The introduction of the Misclassification Likelihood Matrix (MLM) provides a comprehensive view of the model’s misclassification tendencies, enabling decision-makers to identify the most common and critical sources of errors. The results highlight the potential of MLMs in enhancing the interpretability and risk mitigation capabilities of softmax probabilities. The implications of our work extend beyond image classification, with ongoing applications in autonomous systems, such as self-driving cars, to improve the safety and reliability of decision-making in complex, real-world environments. By identifying scenarios where human judgment is preferable to the autonomous system’s assessment, our methodology aims to mitigate the risks associated with distribution shifts and enhance the overall trustworthiness of automated decision-making systems. Future work will focus on several key areas. First, we plan to extend our approach to more complex datasets and network architectures to further validate its effectiveness and generalizability. Second, we aim to investigate the integration of our metric with other techniques, such as uncertainty estimation and domain adaptation, to develop a more comprehensive framework for handling distribution shifts. Third, we will explore the application of our methodology to real-world autonomous systems, such as self-driving cars, to assess its impact on safety and reliability in dynamic environments. Additionally, we plan to conduct user studies to evaluate the interpretability and usefulness of MLMs for decision-makers in various domains. By gathering feedback from domain experts, we can refine our approach to better meet the needs of practitioners and ensure its practical applicability. Finally, we aim to investigate the potential of our approach in other areas, such as medical diagnosis, financial risk assessment, and natural language processing, where the consequences of misclassifications can be severe. By adapting our methodology to these domains, we hope to contribute to the development of more reliable and trustworthy AI systems across a wide range of applications.

### v3

# The Misclassification Likelihood Matrix: Some Classes Are More Likely To Be Misclassified Than Others

@@ 1 Introduction
Additionally, financial systems, where errors might result in significant economic losses; or legal applications, where misjudgments could affect individuals’ lives.
@@
Approaches such as Multi-class Difference in Confidence and Accuracy (MDCA) [hebbalaguppe2022stitch], aim to align CNN-like models’ predicted confidence scores with their actual accuracy across all classes, not just the predicted class (DCA).
@@
Distribution shift is a critical challenge in machine learning that occurs when the statistical properties of the data used to train a model differ from those encountered during deployment or testing [quinonero2009dataset, hendrycks2019benchmarking]. This mismatch can lead to a significant decline in model performance, as the model’s learned patterns may not generalize well to the new data distribution.
Distribution shift can manifest in various forms. Covariate shift occurs when the distribution of input features changes between training and deployment, while the relationship between features and labels remains the same [shimodaira2000improving]. Concept drift refers to situations where the relationship between input features and the target variable changes over time, even if the input distribution remains constant [gama2014survey]. Domain shift encompasses changes in both the input distribution and the relationship between inputs and outputs, often seen when applying a model to a new domain [patel2015visual]. These different types of shifts present unique challenges and require tailored approaches to address them effectively.
To quantify and mitigate distribution shift, researchers have developed a range of sophisticated techniques. Statistical divergence measures, such as Kullback-Leibler divergence [kullback1951information] or Maximum Mean Discrepancy [gretton2012kernel], help quantify the difference between two probability distributions, allowing researchers to detect and measure distribution shift. These measures provide valuable insights into the extent of the shift and can guide mitigation strategies.
Domain adaptation techniques aim to adapt a model trained on one domain (source) to perform well on a different but related domain (target) [wang2018deep, sugiyama2007covariate, pan2009survey, ganin2016domain]. These methods include transfer learning, where knowledge gained from one task is applied to a different but related task, adversarial training, which aims to learn domain-invariant features, and importance weighting, which adjusts the contribution of training samples based on their relevance to the target domain.
Adaptive algorithms are designed to continuously update and adapt to changing data distributions [lu2018learning, baena2006early]. These approaches often involve online learning or incremental learning techniques that can adjust model parameters as new data becomes available. Such algorithms are particularly useful in dynamic environments where the data distribution may evolve over time.
@@ 2.1 Softmax Output And Environmental Aspects
Furthermore, there is a growing awareness of the need for sustainable AI practices and the development of green AI frameworks [schwartz2020green]. These initiatives focus on designing and implementing machine learning systems that prioritize energy efficiency, resource conservation, and environmental sustainability. Sustainable AI practices can help mitigate the environmental impact of discarded softmax outputs and other inefficiencies in machine learning pipelines.
@@ 2.2 Using All Softmax Output Predictions
Utilizing the entire set of softmax prediction probabilities, instead of solely depending on the maximum output, has been widely investigated to improve the safety, robustness, and trustworthiness of machine learning models. By considering the complete distribution of class predictions, more reliable and informative prediction pipelines can be developed, enabling techniques such as uncertainty quantification, anomaly detection, and robustness against adversarial attacks [kendall2017uncertainties, lakshminarayanan2017simple, hendrycks17baseline, goodfellow2014explaining, szegedy2013intriguing].
@@
The Misclassification Likelihood Matrix (MLMs) proposed in this study we believe has potential to further enhance the interpretability and risk mitigation capabilities of softmax probabilities. MLMs provide a comprehensive view of the model’s misclassification tendencies by capturing the likelihood of each class being misclassified as another class. By analyzing the patterns and magnitudes of these misclassification likelihoods, decision-makers can identify the most common and critical sources of error. This information can be used to prioritize model improvements, guide data collection efforts, and establish decision thresholds based on the acceptable level of risk. Moreover, MLMs can be employed to generate explanations for the model’s predictions, highlighting the classes that are most likely to be confused and the factors contributing to the misclassifications. This transparency enables users to better understand the limitations and potential failure modes of the model, promoting informed decision-making and fostering trust in the system.
@@ 3.1 Neural Networks
Figure 1: MNIST testing dataset confusion matrix where the least misclassified digit is 0 (top row) and the most misclassified digit is 8 (second row from bottom up), where the figures represent a snapshot i.e. one single testing dataset, the main difference between a traditional confusion matrix and the MLM which may represent n distribution-shifted testing datasets.
@@
In our analysis, we consider the classifier neural network’s entire output, not just the highest probability score. The network produces a probability vector $\mathbf{p}=(p_{1},p_{2},\dots,p_{K})$, where $\sum p_{i}=1$,
This probability distribution is derived from a logit vector $\mathbf{z}=(z_{1},z_{2},\dots,z_{K})$ through the application of the softmax function:
$p_{i}=\text{softmax}(z_{i})=\frac{e^{z_{i}}}{\sum_{j=1}^{K}e^{z_{j}}}$ (1)
For example, given
$\mathbf{p}=[0.01,0.01,0.01,0.01,0.9,0.01,0.01,0.01,0.01,0.01]$, the predicted class is ’4’, corresponding to the highest value at index five, reflecting the confidence of the prediction for each class from ’0’ to ’9’.
The relationship between logits and probabilities can be expressed as:
$z_{i}=\log\left(\frac{p_{i}}{1-p_{i}}\right)$ (2)
Here, $z_{i}$ represents the logit for class $i$, while $p_{i}$ denotes the probability that the input belongs to class $i$. These logits can be interpreted as log-likelihoods of class membership [goodfellow2016deep, bishop2006pattern].
By examining the entire output vector, rather than just the highest score, we gain deeper insights into the model’s decision-making process and its level of certainty or uncertainty across all possible classes.
@@
The model was trained using the Stochastic Gradient Descent (SGD) optimizer with a learning rate of 0.01, batch size of 64 and Cross-Entropy loss function.
@@ 3.2 Perturbed MNIST Dataset
To test our trained model, we augment by two orders of magnitude the MNIST testing dataset, where the original testing dataset consists of 10,000 images with 10 classes of approximately 1,000 examples each. We create 12 perturbations - Brightness, Contrast, Defocus Blur, Fog, Frost, Gaussian Noise, Impulse Noise, Motion Blur, Pixelation, Shot Noise, Snow and Zoom Blur - following prior work by [hendrycks2019benchmarking], at 10 intensity levels, generating a dataset of 1,210,000 images. We modify perturbation function parameters by trial and error such that increased perturbation levels on the testing dataset cause network predictive accuracies to decrease linearly. The 60,000 image training dataset was left unperturbed. While it would be desirable to use more than one dataset, it has been demonstrated that softmax output distances to centroids (explained in the next section) are greater for misclassified examples in both MNIST/CNN and CIFAR-10/ViT (Vision Transformer) [sikar2024acceptautomatedpredictionsdefer]. CNNs have been the traditional go-to architecture for image recognition tasks, leveraging local receptive fields and spatial hierarchies. Vision Transformers, on the other hand, represent a more recent paradigm shift in computer vision, adapting the self-attention mechanism from natural language processing to image analysis. Despite these fundamental differences in architecture, both types of models exhibit similar behavior in terms of the relationship between softmax output distances and misclassification likelihood.
It is also important to note that the probability distribution of a digit belonging to a class is not normally distributed, but rather follows a multinomial distribution across the 10 digit classes.
@@ 3.3 Misclassification Likelihood Matrix
Figure 2: Perturbed MNIST testing dataset MLM, where dark blue indicates greater, and lighter yellow lower likelihood of misclassification. The figures effectively represent 120 MNIST testing datasets with 10 levels of noise across 12 different perturbation types.
@@
After obtaining the initial centroids, we use them to initialize the K-Means algorithm. The K-Means algorithm iteratively assigns each softmax output to the nearest centroid and updates the centroids based on the assigned outputs until convergence. This process refines the centroids and yields the final adjusted centroids $\hat{\mu}_{0},\hat{\mu}_{1},...,\hat{\mu}_{9}$.
b. Obtaining the Distance to a Neighboring Centroid: Given a test example $x$ with true class label $y$, we obtain the softmax output $s=f_{\theta}(x)$, that we interpret as points in a simplex, which has a natural Euclidean geometry. To calculate the distance between $s$ and a neighboring class centroid $\hat{\mu}c$, where $c\neq y$, we use the Euclidean distance:
@@ 3.4 Testing the hypothesis
Let $L^{(p)}\in\mathbb{R}^{10\times 10}$ be the misclassification likelihood matrix at perturbation level $p$, where $p\in\{1,2,...,10\}$. Each element $L^{(p)}_{y,c}$ represents the likelihood of an example from class $y$ being misclassified as class $c$ at perturbation level $p$.
@@
$\sigma_{y,c}=\sqrt{\frac{1}{10}\sum_{p=1}^{10}(L^{(p)}_{y,c}-\bar{L}_{y,c})^{2}}$ (10)
@@ 4.2 Clustering results
Table 1 presents the Euclidean distances between the average softmax outputs and the K-Means centroids for each digit class (0-9). For most digit classes (0, 1, 2, 3, 4, 7, 9), the Euclidean distances are extremely small, on the order of 1e-14 to 1e-15, indicating that the average softmax outputs are very close to the centroids obtained from the K-Means clustering algorithm. However, for digit classes 5, 6, and 8, the Euclidean distances are slightly larger, on the order of 1e-4, suggesting a minor difference between the average softmax outputs and the K-Means centroids for these specific classes, which could be attributed to the presence of some misclassified or ambiguous examples within these classes. Despite the slight deviations observed for classes 5, 6, and 8, the overall distances between the average softmax outputs and the K-Means centroids are relatively small across all digit classes, indicating a strong similarity between the representations learned by the neural network and the centroids obtained from the K-Means clustering algorithm. The small distances suggest that using the average softmax outputs as initial centroids for the K-Means clustering algorithm is an effective approach, as they provide a good starting point for the clustering process and can potentially lead to faster convergence. Given the data is highly skewed towards the centroids, alternative statistical methods like standard percentiles lack granularity.
@@ 4.3 Misclustered Images
Looking at the pattern of the softmax output (network prediction) indicates how confident a model is of the class prediction. Figure 3 contains two bar charts that show the MNIST-trained CNN softmax output for MNIST dataset digit ids 8688 and 22561 (starting from index 0), where both are labelled as class 6 and correctly classified by the CNN as such. The bar chart on the left shows the softmax output assigns a very high probability to both classes 5 and 6 (where 6 is the highest). The bar chart on the right shows that the CNN assigns a high probability to classes 6 and 8 (where 6 is again the highest). The cluster labels however do not correspond as digit id 8688 is nearer to the cluster 5 centroid, and digit id 22561 is nearer to cluster 8 centroid. This volatility may be explained by the relatively high deltas shown in Table 1, where the K-Means centroids moved a relatively large distance with respect to the initial values (given by average softmax outputs for correct digit classifications in the training dataset), for digit classes 5, 6 and 8.
Conversely, the pattern of the distance to class centroids for a given softmax output indicates how close the softmax output is to a class centroid with respect to other class centroids. Figure 4 shows on the left the distance to all class centroids of digit id 8688 network softmax output, placing it closing to centroids 5 and 6, while the image on the right (digit id 22561) places the softmax output closer to centroids 6 and 9. We notice the inverse relation between both sets of figures, where the highest bars (network confidence) in the first set correspond to the lowest bars (distance to class centroid) set in the second set.
@@ 4.4 Applying Perturbations
Through trial and error, we find sets of parameter values for each perturbation function, such that as the level increases from 1 to 10, the network predictive accuracy decreases from 98.52% maximum to 36.32% minimum and from average 97.61% (level 1) to 48.83% (level 10). The accuracy degradation can be observed both Figures 5 and 6.
@@
Figure 7 shows a heatmap where every cell represents one perturbation (y axis labels, 12 in total) at one intensity level (x axis labels, 10 in total) applied to the 10,000 image MNIST training dataset, the perturbed images are presented to the trained model and the resulting accuracy is placed in the corresponding cell. The heatmap therefore represents an experiment with 1,210,000 predictions on perturbed images of the perturbed dataset as described in section 3.2.
@@ 4.5 Misclassification Likelihood Matrix
Figure 2 shows the Misclassification Likelihood Matrix obtained by the method described in section 3.3. Comparing the figure with the MNIST testing (non-perturbed) dataset shown in Figure 1 we can notice some information added by the MLM that may help inform risk assessment. On the top right of the confusion matrix is a cluster of zeros, representing digits 0, 1 and 2 misclassifications of 8 and 9. The MLM is able to provide fine-grained information about the misclassification likelihood, assigning relatively low albeit non-zero likelihood misclassification values. As the MLM represents a much larger dataset, there are some noticeable discrepancies. The MLM does not assign a high likelihood for digit 9 misclassified as 7, which is the highest in the testing dataset. On the other hand, row 7 is a close match with 7 being misclassified as 2 holding in the perturbed MNIST dataset.
MLM and traditional confusion matrices offer complementary insights into a neural network’s classification performance. The MLM, derived from multiple perturbed versions of the testing set, provides a fine-grained view of misclassification tendencies. It represents the likelihood of each digit class being misclassified as another, with color intensity indicating misclassification probability. In contrast, the confusion matrix shows actual counts of correct and incorrect classifications for a single test set. The MLM offers an additional perspective, showing likelihood values even for rare misclassifications.
While the confusion matrix clearly shows classification performance with higher numbers on the diagonal indicating better accuracy, the MLM requires a different interpretation. Zero values on the diagonal mean one class cannot be misclassified as itself. Both tools highlight patterns like the high likelihood of misclassifying 7 as 2, while the MLM also shows likelihood for corresponding zero values in the confusion matrix e.g. the likelihood of class zero being misclassified as classes 1, 2, 3, 4 or 5.
@@ 5 Conclusions and Future Work
In this study, we proposed a novel approach for quantifying the reliability of neural network predictions under distribution shifts by leveraging clustering techniques and analyzing the distances between softmax outputs and class centroids.
The introduction of the Misclassification Likelihood Matrix (MLM) provides a comprehensive view of the model’s misclassification tendencies, enabling decision-makers to identify the most common and critical sources of errors. The results highlight the potential of MLMs in enhancing the interpretability and risk mitigation capabilities of softmax probabilities.
The implications of our work extend beyond image classification, with ongoing applications in autonomous systems, such as self-driving cars, to improve the safety and reliability of decision-making in complex, real-world dynamic environments. By identifying scenarios where human judgment is preferable to the autonomous system’s assessment, our methodology aims to mitigate the risks associated with distribution shifts and enhance the overall trustworthiness of automated decision-making systems.
Future directions:
1. Incorporate entropy and entropy apex metrics into our framework. The entropy apex, a point approximately equidistant to all class centroids, is expected to be a region of high entropy where misclassified examples cluster.
2. Explore alternative distance metrics, such as cosine similarity, to measure the relationships between class centroids and predictions. This could offer new perspectives on the geometric properties of the feature space and how they relate to classification decisions.
3. Integrate our metric with complementary techniques, including uncertainty estimation and domain adaptation methods. This integration could lead to a more comprehensive framework for handling distribution shifts, combining the strengths of multiple approaches to create more robust and adaptable models.
4. Investigate the application of our methodology to a broader range of real-world autonomous systems, focusing on how it can enhance safety and reliability in diverse and challenging environments.